A 1.5V cell supplies 0.20A to a lamp for 7 hours before the lamp goes out. What is a sensible estimate for the initial chemical energy content of the cell? A) 1*10^2J B) 1*10^4J C) 1*10^6J D)1*10^8J


1
Expert's answer
2019-10-02T09:38:30-0400

Total energy given to lamp = initial chemical energy of of cell= "Vit"

Where,

"V= voltage=1.5 \ V\\i=current=0.2\ A\\t=time=7\times 3600 \ sec"

Total initial chemical energy = "1.5\times 0.2\times 7\times 3600=7560 \ J"

7560 is of the order of "10^4"

So sensisble estimate of initial chemical energy="1\times10^4\ J"

So option (B) is correct